### v1

# JWST Observations of the Double Nucleus in NGC 4486B: Possible Evidence for a Recent Binary SMBH Merger and Recoil

Abstract
A recent study of the compact elliptical galaxy NGC 4486B using JWST–NIRSpec IFU kinematics confirmed a supermassive black hole (SMBH) of mass ${\cal M}_{BH}=3.6\pm 0.7\times 10^{8}$ ($\sim 8$% of the stellar mass). In addition to its double nucleus, the nuclear kinematics show pronounced asymmetries: a velocity-dispersion peak displaced by 6 pc from the galaxy center and a $\sim 16\,\mathrm{km\ s}^{-1}$ offset in the mean stellar line-of-sight velocity near the SMBH. We examine the origin of the 12 pc double nucleus and these asymmetries and show that the observations favor an SMBH surrounded by an eccentric nuclear disk (END). END formation models require the SMBH to experience a GW recoil following a binary SMBH merger. Our orbit-superposition models contain $\sim 50\%$ retrograde stars at the edge of the nuclear region, in striking agreement with END-formation simulations. We infer a pre-merger mass ratio $q>0.15$ and a recoil kick of $\sim 340\,\mathrm{km\ s}^{-1}$. Our $N$-body simulations show that with such a kick, the SMBH returns to the center within $\sim 30$ Myr. Its flat central core is also consistent with earlier “binary black hole scouring”. We test two alternative mechanisms—buoyancy-driven oscillations and a pre-merger SMBH binary—but neither reproduces the observed offsets, favoring the GW-kick scenario. Our direct $N$-body simulations further show that a prograde SMBH binary in a rotating host can stall in a corotation resonance, delaying coalescence. Thus, although NGC 4486B is an old, relaxed galaxy near the Virgo cluster center, its SMBH appears to have merged only recently, making its nucleus a rare nearby laboratory for studying post-merger SMBH dynamics.
Stellar Dynamics (1596) — Supermassive black holes (1663) — Ultracompact dwarf galaxies (1734) — Galaxy evolution(594) — Compact galaxies(285) —
1 Introduction
NGC 4486B (VCC1279) is a cE galaxy at the center of the Virgo Cluster, projected $7.3^{\prime}$ ($\sim 35$ kpc) from the massive elliptical M87, with a modest line-of-sight velocity difference of $\sim 200\,\mathrm{km}\,\mathrm{s^{-1}}$ that indicates a close dynamical association. NGC 4486B is characterized by a small effective radius of $2.33^{\prime\prime}$ ($\sim 0.19$ kpc) and an absolute magnitude of $M_{V}\sim 17.6$, with a total stellar mass of $_{*}\sim 6\times 10^{9}MM_{\odot}$ (Lauer.1996; Ferrarese.2006; Janz.2016). Recently, we showed using JWST-NIRSpec IFU kinematical data and archival HST photometry, that NGC 4486B hosts a supermassive black hole (SMBH) with a mass of ${\cal M}_{BH}=(3.6\pm 0.7)\times 10^{8}\mbox{$M_{\odot}$}$, making this black hole significantly overmassive (${\cal M}_{BH}/M_{*}\sim 8$%) relative to expectations from black hole scaling relations (Behzad.2025, hereafter Paper 1). This suggests that NGC 4486B is an extreme case of a tidally stripped galaxy, likely the compact core of a once significantly more massive system.
[…]
One plausible explanation for the double nucleus in M31 and NGC4486B is the presence of an eccentric, apsidally aligned stellar disk orbiting a central SMBH (Tremaine.1995; Peiris.2003; Bacon.2001; Sambhus.2002; BrownMagorrian.2013; Wernke.2021). In this model, the nearly Keplerian potential of the SMBH dominates the stellar dynamics, shaping a lopsided configuration where stars follow coherent, eccentric orbits. The apparent double nucleus arises from the asymmetric distribution of stellar density along these orbits: the brighter component corresponds to stars near their apocenters, where they move more slowly, while the fainter component is associated with stars near their pericenters, close to the black hole.
Lauer.1996 explored a range of other possible origins for the double nucleus in NGC 4486B. A dust absorption feature as well as an infalling globular cluster were ruled out due to the nearly identical morphologies in $V$ and $I$ bands and the lack of significant color gradients across the nucleus. An ongoing merging or binary stellar nucleus was also deemed unlikely, as the decay timescale for such a system due to dynamical friction is short ($\lesssim 10^{7}$ yr), and there is no evidence for a recent merger in NGC 4486B. Projection effects from unrelated foreground or background sources were likewise dismissed based on the nearly identical colors and recession velocities of the two nuclei (Kormendy.1997).
A striking feature of NGC4486B is its flat central density core (see Section 2.1). One of the best accepted mechanisms for producing such cores in early type galaxies is via “binary-black-hole scouring” following the merger of two gas poor galaxies, each with its own SMBH. The black holes (BHs) quickly sink to the common center of the galaxy via dynamical friction and become bound with separations of 10s of parsecs. In order to shrink the orbit further, the SMBH binary transfers its orbital kinetic energy to surrounding stars, via the $3$-body slingshot effect, preferentially ejecting stars with low angular momentum (e.g. Milos.2003; Baile15) until the binary separation becomes small enough for gravitational wave (GW) radiation to dominate. High-resolution $N$-body simulations (khan+12a; Rantala2024) have shown that an SMBH binary can eject a stellar mass amounting to 252-5 times its total mass from the central region, thereby converting a central cusp of stars into a flatter central core (rantala2018; Nasim2021; Khan2025).
Numerical relativity simulations (e.g., Campanelli.2007ApJ; Campanelli.2007PRL) have shown that the anisotropic emission of GWs during the merger of two SMBH delivers a kick to the final black hole with a magnitude that depends on a variety of factors: the mass ratio $q=m_{2}/M_{1}$ of the secondary to the primary; whether the BHs have spin; the alignment of their spins to their mutual orbital plane; and whether the black hole spins are aligned or misaligned. While kicks as large as $\sim 2000-4000\,\mathrm{km\ s}^{-1}$ (Campanelli.2007PRL) are possible, the actual distribution of kick velocities is uncertain and most kicks are probably $\lesssim 1000$$\mathrm{km\ s}^{-1}$(Gualandris.2008). Even a modest kick to the SMBH causes it to oscillate transfering orbital energy to nearby stars, further flattening the central density profile (Gualandris.2008; Khonji.2024).
Recently, the eccentric nuclear disk (END) model has received renewed interest due to a suite of direct $N$-body simulations of nuclear disks surrounding SMBHs (Akiba.2021; Akiba.2023; Bright.2024). These authors demonstrate that a GW recoil kick can naturally transform a circular stellar disk bound to an SMBH into an eccentric, apsidally aligned lopsided one. In this scenario, the stars that remain bound to the SMBH after the kick experience angular-momentum changes that place them on eccentric orbits, whose net alignment is oriented roughly orthogonal to the kick direction. For a sufficiently large kick, stars in the outer parts of the disk can even become retrograde in the rest frame of the SMBH.
In this work, we investigate the photometric and kinematic signatures of the double nucleus in NGC 4486B and build upon our previous dynamical modeling results from Paper 1, as well as previous $N$-body simulations, to examine the origin and dynamical state of its END. Using a self-consistent $N$-body realization constructed using our Schwarzschild orbit-superposition model from Paper 1, we perform simulations exploring both pre- and post-merger scenarios. We use previous analysis from Akiba.2021; Akiba.2023; Akiba.2025 and the observed properties of the double nucleus to estimate the kick velocity and likely mass-ratio of the black hole binary. We simulate various recoil kicks to assess the timescale on which the black hole would sink to the center. In addition to GW recoil, we test other alternative mechanisms capable of displacing the SMBH and influencing the observed kinematics, such as buoyancy-driven displacements of the SMBH (Cole2012; Banik2021; Banik2022; Dattathri2025a), and the pre-merger evolution of a prograde binary black hole merger (holley+15; Rasskazov+16; Mirza+17; Khan2020). We also quantify the expected amplitude of SMBH Brownian motion (Merritt.2001) to verify that the displacements seen in our simulations and inferred from observations exceed the stochastic wandering expected from finite stellar encounters Brownian motion.
While nearly two decades of theoretical and simulation work have explored the consequences of binary black hole mergers, the resulting GW recoils, and SMBH Brownian motion in galaxies, NGC 4486B appears to be the first system exhibiting multiple observable signatures of a recent merger—making it a benchmark case for studying the dynamical aftermath of SMBH coalescence. Although NGC 4486B and M31 are among the best-studied examples, double nuclei are not unique. Debattista.2006 examined a sample of dE galaxies from Lotz.2004 imaged in the WFPC2 snapshot surveys GO-8600 and GO-6352 (PI: H. Ferguson) and found that VCC 1107, FCC 208, and VCC 128 show signs of double nuclei, with VCC 128 exhibiting two comparably bright components and the closest separation ($\sim 32$ pc). Tracing the formation and evolution of the eccentric disk in NGC 4486B and other galaxies may provide indirect constraints on the frequency of SMBH mergers – an essential parameter for interpreting the stochastic GW background measured by Pulsar Timing Arrays (PTAs) (Nanograv.2023; nanograv.2024; EPTA.2023; EPTA.2024).
The structure of this paper is as follows. In Section 2, we describe the HST and JWST datasets used in our analysis and briefly summarize the Schwarzschild dynamical modeling from Paper 1. In Section 3, we present the observational evidence for an END in NGC 4486B and compare the nuclear structure and kinematics of our Schwarzschild model with simulations of ENDs formed through GW recoil kicks. Using the observed properties of the double nucleus, we derive constraints on the black hole binary mass ratio and the required recoil velocity. In Section 4, we use direct $N$-body simulations to follow the post-kick trajectory of the SMBH and evaluate the timescale on which it returns to the galaxy center, as well as the kinematic signatures such a kick would imprint. Section 5 examines alternative mechanisms that could produce an off-center SMBH or similar kinematic asymmetries. Finally, in Section 6 we summarize our findings and discuss their implications. Throughout this study, we adopt a distance of $16.3$ Mpc to NGC 4486B (Blakeslee.2009)
2 Observational data and overview of Schwarzschild models in Paper 1
In this section we briefly summarize the properties of photometric imaging from Hubble Space Telescope (HST) instruments WFPC2 and ACS/WFC and the spectroscopic data from JWST-NIRSpec IFU that were used to construct the Schwarzschild dynamical models of NGC 4486B presented in Paper 1 (Behzad.2025). In addition, we briefly summarize the main results from that paper including details about the best-fit Schwarzschild models that are used in this work.
2.1 Photometric data and image analysis
model $n$ $r_{e}$ (arcsec) $\Sigma_{e}$ $\mathrm{mag/arcsec^{2}}$ $q$ PA (degree) $\chi^{2}/N_{\mathrm{dof}}$ AIC
1 Sérsic 3.05 2.78 18.58 0.79 88.67 5.5 30032248
2 Sérsic 1.2 0.78 16.45 0.65 87.90 0.67 365042
1.80 2.93 18.72 0.86 89.29
3 Sérsic 1.26 0.40 16.73 0.74 90.55 0.42 227140
1.76 1.33 18.01 0.71 86.86
1.46 3.26 19.08 0.90 91.47
4 Sérsic 1.44 0.39 17.27 0.74 100.63 0.36 192528
1.39 0.52 17.44 0.73 84.85
2.30 1.80 19.47 0.57 85.24
1.58 3.02 19.04 0.93 96.21
Table 1: The GALFIT output for the best-fit 2D surface brightness models of NGC 4486B F850LP image using one, two, three, and four Sérsic functions. Columns from left to right list the number of Sérsic components, Sérsic indices ($n$), effective radius ($r_{e}$), surface brightness at the effective radius ($\Sigma_{e}$), flattening ($q$), position angle (PA), reduced chi-square ($\chi^{2}/N_{\mathrm{dof}}$, where $N_{\mathrm{dof}}$ is the number of degrees of freedom), and the Akaike Information Criterion (AIC), defined as $\chi^{2}+2k$, with $k$ representing the number of model parameters for each model.
In Paper 1, we used archival HST-ACS imaging in the F850LP filter (pixel scale $\sim 0.05\arcsec$ $\lesssim 4$ pc and PSF FWHM $\sim 0.9\arcsec$ $\lesssim 72$ pc) acquired as part of the ACS Virgo Cluster Survey (Cote.2004). A model point spread function (PSF) was generated using the TinyTim software package (Krist.2011) and used for all photometric analyses. The full photometric frame reaches out to $\sim$1 kpc in radius, with an effective radius of $2.33\arcsec$ (182 pc), extending well beyond the field of view of the NIRSpec-IFU observations. We also examine the WFPC2 F [^5] W image from Lauer.1996, which provides slightly higher spatial resolution in the central region and reveals the double nucleus structure more distinctly.
Previous analyses of the morphology, isophotal parameters, and surface brightness profile of NGC 4486B showed that it exhibits a shallow surface brightness profile in its central region (Lauer.1996; Faber.1997; Ferrarese.2006). Ferrarese.2006 modeled its surface brightness using a single Sérsic component and obtained $n=2.73$ and $r_{e}=2.33\arcsec$. We repeated the 2D surface brightness fitting of the ACS F850LP image using GALFIT (GALFIT2010) models with one, two, three, and four Sérsic components. For a single component, the Sérsic index is $n=3.05$, slightly higher than in previous work. As the number of Sérsic components increases to three, the quality of the fit improved. Beyond this point, however, the model did not show substantial improvement (see Table 1). By adding more components, the Sésic index of the inner region decreases to a range of n=1.21.4n=1.2-1.4. As might be expected due to the double nuclei, the inner region of NGC 4486B cannot be adequately described by single or multiple nested Sérsic components (even allowing for the components to have different centers). The major-axis position angles of all components across different models are similar and consistent with the position angle profile derived from the ellipse fitting to the isophotes. The ellipticity and position angle of the isophotes (outside the innermost $\sim 0.15\arcsec$ region occupied by the double nucleus) remain relatively constant with radius, indicating that the galaxy is close to axisymmetric overall. Similar results were obtained with the WFPC2 F [^5] W image.
Figure 1: The top panel presents the projected luminosity density of NGC 4486B (black dots) derived from the ACS F850LP image along with various models (see text for details). The bottom panel displays the deprojected 3D luminosity density profiles along the semi-major axis. The vertical dashed line indicates the PSF FWHM of the ACS F850LP image.
Figure 2: First row, right panel: Unsharp-masked WFPC2/F [^5] W image of the double nucleus in NGC 4486B, showing two peaks separated by $\sim$12 pc. First row, left and middle panels: Stellar kinematic maps of $v$ and $\sigma$ from JWST-NIRSpec IFU data. In all panels, surface-brightness contours from the deconvolved WFPC2/F [^5] W image are overplotted. The black dot marks the center of the faint peak (P2), which coincides with the $\sigma$ peak from NIRSpec and is identified as the likely location of the SMBH. The purple dot indicates the center of the bright peak (P1), coinciding with the center of the IFU data cube. The vertical dashed line marks the kinematic center. Second row: One-dimensional kinematic profiles of NGC 4486B derived from JWST-NIRSpec IFU data, showing the $|v|$, $\sigma$, and $|v|/\sigma$ values along the major axis intersecting both nuclei. Red points represent measurements and associated uncertainties from the right side of the kinematic maps, where $\sigma$ peaks. Blue points correspond to the left side of the galaxy, mirrored about the rotation axis for direct comparison. The gray shaded region marks the approximate radial extent of the eccentric disk inferred from photometric and kinematic features (corresponds to the white rectangle above).
In Paper 1, we used the multi-Gaussian expansion (MGE) method (Cappellari.2002) to model the surface brightness of NGC 4486B. The details of the fit are provided in Table 1 of Paper 1. In the top panel of Figure 1, we present the projected luminosity density of NGC 4486B (black dots), along with the corresponding models: the MGE model (in red) and Sérsic models with one (green), two (blue), three (orange), and four (gray) components. The vertical dashed line indicates the PSF FWHM of the F850LP image. None of the models with multiple Sérsic components are able to replicate the flatness of the inner profile as well as the MGE model does in this region. To test whether the presence of the double nucleus affects the derived surface brightness profile, we allowed the centers of the Sérsic (or Gaussian) components to vary freely in our GALFIT modeling. In the best-fit model, the centers of all components are found to lie within $\leq 1$ pixel of each other. Likewise, in the standard MGE fitting, where all Gaussian components share a fixed center at the brightest peak, the result remains consistent. The choice of fixing or freeing the component centers therefore does not significantly affect the shape of the final luminosity profile. This is likely because, although the nuclear region is slightly lopsided rather than fully symmetric, the two peaks are separated by only $\simeq$ 3-4 pixels, comparable to the PSF FWHM. Consequently, the central light distribution appears smooth rather than distinctly double in the F850LP image, causing all model components to converge to nearly the same photocenter, which lies on the bright peak.
For the MGE model we determined the 3D luminosity density of NGC 4486B from the projected luminosity density profiles by deprojecting the 2D MGE components into a 3D axisymmetric distribution, assuming an inclination angle of $\theta=70^{\circ}$ (inferred from the Schwarzschild modeling in Paper 1). For the Sérsic models, we numerically compute the 3D luminosity density by calculating the deprojection integral, assuming spherical geometry. The bottom panel of Figure 1 displays the deprojected 3D luminosity density profiles of each model along the semi-major axis. The inner luminosity profile of the MGE model is shallower than the multi-component Sérsic models, which aligns with expectations from the 2D surface brightness profiles.
The luminosity density profile in Figure 1 (lower panel) shows that NGC 4486B has a core that extends out to $\sim 0.25\arcsec$ (20 parsecs). The luminosity density $\nu(r)$ in this region is described by $\nu(r)=r^{-\gamma}$ where the cusp slope $\gamma=0.14$.
As mentioned in Section 1, the favored mechanism for core formation in elliptical galaxies is “core scouring by an SMBH binary”. The low-density core produced in this manner can extend up to five times the black hole binary influence radius (rantala2018; Nasim2021; Khan2025). For the black hole in NGC 4486B ${\cal M}_{BH}=3.6\times 10^{8}$$M_{\odot}$ (Paper 1), the sphere of influence is estimated to be 203520-35 pc, which is comparable to the observed core radius.
The surface brightness fitting above ignored the double nucleus, which is barely resolved in the ACS/WFC F850LP image. In the higher spatial resolution of the HST WFC2 F [^5] W image (Lauer.1996) however, it is clearly resolved (Figure 1 of Paper 1 gives both images). Figure 2 (first row, right panel) shows an unsharp-masked image of NGC 4486B, in which the double nucleus is revealed with striking clarity. The two peaks are distinctly separated by $\sim$12 pc, with the fainter component (P2) marked by a black dot while the brighter nucleus (P1) is marked by a purple dot. Photometrically, P1 has a central surface brightness of $\mu_{I}=12.89$ mag arcsec-2, while P2 has $\mu_{I}=13.00$ mag arcsec-2, with a minimum of $\mu_{I}=13.06$ mag arcsec-2 between them (Lauer.1996). The white rectangle in this image encloses the nuclear region that we consider in greater detail below.
2.2 JWST NIRSpec-IFU kinematics
[…]
The kinematic datacube was Voronoi-tesselated (Cappellari.2003_voronoi) to achieve a uniform signal/noise ratio of $\sim 80$ across the NIRSpec field-of-view. Line-of-sight stellar kinematics were extracted using the penalized pixel-fitting method (pPXF; Cappellari.2023), fitting each Voronoi-binned spectrum with synthetic templates from the PHOENIX library (Husser.2013). We model the line-of-sight velocity distribution (LOSVD) using a Gauss-Hermite (GH) expansion as implemented in pPXF to extract mean velocity $v$, the width $\sigma$ of the best-fit Gaussian, and deviations captured by $h_{3}$ (related to skewness) and $h_{4}$ (related to kurtosis) (e.g., Marel.1993; Gerhard.1993).
The left and middle panels of the first row in Figure 2 show $v_{\mathrm{los}}$ (in the galaxy rest-frame) and the velocity dispersion maps, respectively, obtained from JWST-NIRSpec. The white rectangle encloses the same spatial region in all three panels. The left-hand panel shows that the galaxy’s central velocity (vertical dashed line) lies between P2 (black dot) and P1 (purple dot), but is closer to P1. It is clear from the middle panel that the peak of the velocity dispersion (reddest bin) coincides with the fainter peak (black dot), which we therefore assume to be the position of the SMBH. The $\sigma$ peak, measured from NIRSpec, has a maximum value of $\sigma=287.2\pm 7.4\,\mathrm{km\ s}^{-1}$. This offset in the velocity dispersion peak is also seen in the second row of Figure 2, which shows 1D-kinematic profiles through the major axis of NGC 4486B. In this figure, the left side of the kinematic maps is mirror-reflected about the rotation axis to enable easier comparison of the two sides, and $|v|$ and $\sigma$ along the major axis are plotted. The vertical grey band shows the region encompassed by the white rectangle in Figure 2. In addition to the fact that the velocity dispersion profile is not symmetric about the center, we see in the top panel that $|v|$ of stars in the vicinity of the presumed SMBH is higher than for stars on the opposite side of the galaxy by $16\,\mathrm{km\ s}^{-1}$ The lowest panel shows the ratio $|v|/\sigma$, indicating that this galaxy has roughly equal amounts of rotational and dispersion support over the range of radii examined.
2.3 Schwarzschild model and $N$-body model for NGC 4486B
To constrain the large-scale mass distribution, in Paper 1 we supplemented the JWST data with long-slit kinematics from CFHT/SIS (Kormendy.1997), using only the outer regions ($r\gtrsim 0.9\arcsec$) to minimize resolution-related discrepancies. The two-dimensional kinematic maps and the unsharp-masked image in Figure 2 reveal a well-resolved double nucleus and a velocity-dispersion peak offset from the photometric center. In Paper 1, we combined these long-slit kinematics with the photometric data described in Section 2.1 and the JWST kinematics (Section 2.2) to construct dynamical models of the galaxy and determine the SMBH mass. We applied two independent dynamical approaches: axisymmetric Schwarzschild orbit-superposition modeling (e.g. Schwarzschild.1979; vanderMarel.1998; Cretton.1999; Gebhardt.2003; Valluri.2004; Thomas.2004; vandenBosch.2010; Vasiliev.2020A) (which fits the full LOSVD including higher-order moments), and axisymmetric Jeans Anisotropic Modeling (JAM, Cappellari.2008) (which used only the first two velocity moments). We constrained the black hole mass as well as other properties of NGC 4486B. The two methods gave us broadly similar results and we refer readers to Paper 1 for further details. We restrict our discussion in this work to the orbit superposition models that were constructed with the FORSTAND code (Vasiliev.2020A).
As both JAM and Schwarzschild modeling assume global symmetry and steady-state equilibrium, they cannot reproduce the pronounced asymmetries in the JWST-NIRSpec IFU kinematics and light distribution associated with the double nucleus. To gauge the impact of this modeling limitation, in Paper 1 we analyze three data configurations: (i) the original asymmetric kinematic maps; (ii) maps in which spaxels at and around the peak of the velocity dispersion, $\sigma$, are masked across all moments, suppressing the strongest asymmetry allowing us to probe the minimum black hole mass allowed; and (iii) a final test in which only the $\sigma$ map was modified by shifting its peak to the photometric center, an intentionally unphysical but informative symmetrization used to explore the maximum mass permitted under the symmetry assumption. These models yielded $2.8^{+0.6}_{-0.4}\times 10^{8}\leq{\cal M}_{BH}/\mbox{$M_{\odot}$}\leq 5.3^{+0.9}_{-0.9}\times 10^{8}$ with the detection significant at $3\sigma$ in all models. The inclusion or exclusion of a dark matter halo has a negligible impact on the measured ${\cal M}_{BH}$. Our Schwarzschild models, fit directly to the original JWST kinematics (no shifting or masking of the peak), yielded a black hole mass ${\cal M}_{BH}=3.6\pm{0.7}\times 10^{8}$$M_{\odot}$, a stellar mass-to-light ratio of $M/L_{[F850LP]}=3.9\pm 0.5$ corresponding to a stellar mass $M_{*}=7.0\pm 0.8\times 10^{9}\mbox{$M_{\odot}$}$, and an inclination $\theta\sim 70^{\circ}$.
In addition to providing the globally optimal model parameters, the Schwarzschild method delivers the full orbital distribution function through the weights assigned to each orbit in the library. These weights allow us to generate a self-consistent, data-driven $N$-body realization of the galaxy. Although one of our best-fitting models includes a modest dark matter fraction ($\sim 14\%$ by mass), for the purposes of this study, we construct the $N$-body model using only the stellar component in order to simplify the subsequent dynamical experiments. The initial conditions for our simulations are drawn from the Schwarzschild model with ${\cal M}_{BH}=3.51\times 10^{8}\mbox{$M_{\odot}$}$, and total stellar mass $M_{*}=7.44\times 10^{9}\mbox{$M_{\odot}$}$, and an intrinsic axisymmetric shape with axis ratio $0.83$.
3 Evidence of an Eccentric Disk formed around a recoiling SMBH
3.1 Comparison of nuclear structure and kinematics of NGC 4486B with eccentric nuclear disk simulations
Wernke.2021 used $N$-body simulations to produce mock photometric and two-dimensional kinematic maps of ENDs for a wide range of viewing orientations. Their results show that the observed morphology depends strongly on the angle between the eccentricity vector (the disk’s major axis) and the line of sight. When the disk is viewed edge-on with its major axis aligned along the line of sight, only a single, centered nucleus is seen. If the disk is viewed edge-on with its minor axis along the line of sight, the nucleus appears offset. For edge-on or nearly edge-on views at intermediate angles relative to the eccentricity vector, the system appears as a double nucleus, where the fainter secondary peak lies close to the location of the black hole (see Figure 2 in Wernke.2021). They also found that the velocity dispersion is generally elevated and peaks at the location of the SMBH, irrespective of the viewing angle, and this dispersion peak typically does not coincide with the photometric center (brightest peak) unless the END is viewed such that it appears as a single centered nucleus.
We therefore expect a double nucleus in NGC 4486B, consistent with our nearly edge-on viewing angle ($\theta\sim 70^{\circ}$), inferred from our Schwarzschild models. The morphology in the right-hand panel of Figure 2 closely resembles the mock ENDs viewed nearly edge-on at intermediate angles to the eccentricity vector in Wernke.2021, supporting this interpretation and consistent with the earlier conclusion of Lauer.1996. In such nearly edge-on ($\theta\sim 70^{\circ}$) configurations, the velocity-dispersion peak is expected to lie near the SMBH rather than at the brightest photometric peak—a behavior clearly seen in Figure 2 (first row, middle). A further examination of the 1D kinematic profiles of NGC 4486B along the major axis intersecting both nuclei (the second row of Figure 2) shows a clear asymmetry between the left and right sides of the galaxy: $|v|$ is elevated by $\sim 16\,\mathrm{km\ s}^{-1}$ and $\sigma$ is elevated by $\sim 57\,\mathrm{km\ s}^{-1}$ compared to the opposite side of the galaxy. Similar kinematic asymmetries and the offset of the velocity dispersion peak from the brighter peak strongly support the eccentric disk simulations (Wernke.2021), which also suggests local dynamical disturbances near the velocity dispersion peak (the presumed location of the SMBH). The enhanced $|v|$ may reflect streaming motions within the eccentric disk or, alternatively, residual motion of the SMBH relative to the center of the galaxy. Since the mock kinematics and images in Wernke.2021 were produced from simulations of isolated ENDs surrounding SMBH, whereas the one in NGC 4486B is embedded in an elliptical galaxy, additional simulations are needed to fully understand the offsets.
Figure 3: Top: Radial profile of the retrograde-orbit fraction from the best-fit Schwarzschild model using the high-resolution orbit library (100,000 orbits). The vertical dashed line at $0.25\arcsec\simeq 20$ pc marks the boundary of the eccentric-disk region highlighted by the white rectangle in Fig. 2. Bottom: Distribution of orbits in phase space, showing mean radius versus orbital circularity. The size of each circular marker scales with the relative weight of the corresponding orbit.
As mentioned in Section 1, one of the proposed formation pathways for ENDs is a GW recoil of an SMBH following coalescence (Akiba.2021; Bright.2024). Such a recoil can drive stars that were initially on nearly circular prograde orbits onto eccentric trajectories and flip a substantial fraction, particularly at larger radii, into retrograde motion, as the displaced SMBH effectively outruns the outer stars and reverses their angular momenta in its rest frame.
The resulting disk therefore contains a mixture of prograde and retrograde orbits, with the retrograde fraction increasing with radius; more powerful kicks generate a correspondingly larger retrograde population (Bright.2024). Simulations further indicate that this transformation occurs on very short timescales - within only a few dynamical times, $\sim 10^{4}-10^{5}$ yr on parsec scales (Akiba.2021). In cases where the disk hosts a substantial retrograde component, apsidal precession slows dramatically, allowing the lopsided eccentric disk structure to survive for $\gtrsim 10^{8}-10^{9}$ yr as a long-lived dynamical fossil of the recoil (Madigan.2018).
Although our Schwarzschild dynamical models from Paper 1 do not explicitly model the double nucleus, the fact that they were designed to fit the observed kinematics provides a window into the orbital architecture of the nuclear region. We now examine the best-fit Schwarzschild model from Paper 1 to assess how the fraction of retrograde orbits varies with radius. For each orbit we extract mean radius $R$, and orbital circularity $\lambda=L_{z}/L_{\rm circ}$ (retrograde if $\lambda<0$) ($L_{z}$ is the angular momentum of the orbit and $L_{\rm circ}$ the angular momentum of a circular orbit with the same binding energy. We keep the orbits with $0\leq R\leq 1^{\prime\prime}$, then divide this range into 20 uniform bins. Since the orbit weights ($w$) in the Schwarzschild model correspond to the mass of stars on a given orbit, in each bin we compute the weighted share of retrograde orbits (retrograde weight divided by total weight in that bin). We compute uncertainties that reflect both the retrograde fraction and the number of effectively independent orbits contributing to a bin, reporting central 68% credible intervals using a Jeffreys prior binomial model (Jeffreys.1946) with an effective sample size derived from the orbit weights. To test robustness against discreteness noise and finite library size, we repeated the modeling in Paper 1 and measurement of the retrograde orbit fraction with an orbit library five times larger (100,000 orbits) than our standard model in that paper. The resulting profiles from the original and large-library model agree within the quoted uncertainties, indicating that our conclusions are not driven by sampling effects. We only show the results for the large-library model. The top panel of Figure 3 shows the retrograde-orbit fraction as a function of radius for the Schwarzschild model, while the bottom panel shows the corresponding orbital circularity distribution. The vertical dashed line at $R=0.25^{\prime\prime}$ ($\sim 20\,\mathrm{pc}$) marks the edge of the eccentric disk region in the white rectangle in Figure 2. Our results closely resemble the orbital distribution of an eccentric nuclear disk formed via a recoil kick (Akiba.2023), which also exhibits a substantial retrograde population ($\sim$50%), particularly in the outer regions of the disk. In Figure 14 of Akiba.2023, the top panel shows the orbital circularity distribution and the bottom panel shows the retrograde-orbit fraction. Although our Schwarzschild models assume axisymmetry and therefore cannot reproduce the full asymmetric structure of the eccentric disk, the emergence of a substantial retrograde component is an inevitable outcome of reproducing the observed JWST kinematics.
Cretton.1999 demonstrated that Schwarzschild models can sometimes assign artificially large weights to retrograde orbits when the kinematics are constrained using only the GH expansion of the LOSVD. In their study of a rapidly rotating S0 galaxy, models without a central black hole reproduced the measured $h_{3}$ and $h_{4}$ moments by populating a large fraction of retrograde (counter-rotating) orbits, even though the observed velocity profiles showed no such component. This “spurious counter-rotation” arises because truncated GH series cannot uniquely describe the full LOSVD, allowing the model to exploit retrograde orbits as a fitting freedom. Importantly, this effect is most pronounced in cold, rotationally supported systems with $|v|/\sigma\gtrsim 1.5$, where small distortions in the LOSVD skewness can be mimicked by a retrograde tail.
By contrast, NGC 4486B is a much more dispersion-dominated system: as shown in the bottom right panel of Figure 2, we measure $|v|/\sigma\lesssim 0.45$ in the inner regions, increasing only to $\sim 0.6$ at the outer edge of our JWST-NIRSpec field of view. In this regime, the degeneracies highlighted by Cretton.1999 are much less severe, making it unlikely that the strong retrograde signal we find is purely an artifact of the GH parameterization. Instead, the retrograde component appears to be an intrinsic dynamical requirement of the system.
To test the importance of retrograde orbits explicitly, we re-ran our best-fit Schwarzschild models with identical initial conditions but excluded all retrograde orbits from the library. This increased the $\chi^{2}$ of the best-fit model by $\sim 23\%$, and the $\sigma$ map becomes vertically elongated and flattened along the projected minor axis. In contrast, the model including retrograde orbits produces a more rounded $\sigma$ structure, in much closer agreement with the observed data. Figure A1 in the Appendix compares the resulting two-dimensional $v$ and $\sigma$ maps: the first row shows the best-fit model with retrograde orbits, while the second row shows the best-fit model with all retrograde orbits removed. This confirms that the retrograde component is required to reproduce the kinematics of NGC 4486B. The presence of a significant retrograde orbit fraction in the region that appears to be dominated by the END thus offers tantalizing evidence that the disk formed as a result of a past SMBH recoil event.
In addition to the recoiling black hole scenario, simulations show that the hardening of an SMBH binary can itself produce a large retrograde fraction in the nuclear orbital distribution: three-body slingshot interactions preferentially deplete low-$L$ and co-rotating (prograde) stars, leaving a tangentially biased remnant in which stable, counter-rotating orbits dominate inside the binary’s sphere of influence (e.g., Meiron.2010; Rantala.2019). Given that NGC 4486B hosts both a flat stellar core and a central SMBH, a binary-scouring scenario provides an alternative explanation for the large retrograde fraction we infer from the Schwarzschild model, offering complementary support for the possibility that the black hole in NGC 4486B is the result of the merger of a binary black hole.
3.2 Constraints on the Recoil Kick and SMBH Binary Parameters
Figure 4: Expected GW recoil kick magnitude, $V_{\rm{kick}}$, as a function of the pre-merger black hole mass ratio, $q$. The thick purple line shows the mean, and the shaded region a $1\sigma$ confidence interval estimated via Monte Carlo simulations. The recoil kick required to induce the double nucleus of NGC 4486B is shown in the gray horizontal dashed line, and the mass of the corresponding pre-merger companion to NGC 4486B’s central SMBH is indicated on the top horizontal axis. The gray vertical dashed line marks $q=0.15$, the lower-limit pre-merger black hole mass ratio for NGC 4486B. Adapted from Akiba.2025 which uses the analytical prescription from Lousto.2010; Lousto.2012.
Akiba.2021 simulated disks with stars initially in circular orbits about an SMBH. Following an in-plane GW recoil kick imparted on the black hole, the stars preferentially align their eccentricity vectors perpendicular to the recoil kick with magnitude
$|\vec{e}_{\rm{avg}}|=\frac{3}{2}\,\frac{V_{\rm{kick}}}{v_{c}}\,$, (1)
where $v_{\rm{kick}}$ is the recoil kick speed and $v_{c}$ is the initial average circular speed of the stars. In their simulations, they found that immediately after the kick, the mean eccentricity jumps to $|\vec{e}_{\rm{avg}}|\sim 1$ but subsequently decreases via large amplitude oscillations to smaller values. Simulations show that while the disk gradually diffuses in semi-major axis and orientation (as $|\vec{e}_{\rm{avg}}|$ decreases), $R_{c}$, the characteristic radius of the disk and the degree of apsidal alignment remain stable over $\gtrsim 10$ Myr (Akiba.2021; Akiba.2023). Akiba.2023 noticed that by setting $|\vec{e}_{\rm{avg}}|\,\sim 1$, one can determine $R_{c}$ of the resulting eccentric disk as,
$R_{c}\equiv\frac{4}{9}\,\frac{GM_{\rm BH}}{V_{\rm{kick}}^{2}}\,$, (2)
where $M_{\rm BH}$ is the mass of the recoiling black hole. This $R_{c}$ is where we expect the eccentricity and apsidal alignment of the post-kick disk to be maximized. Applying these analytics to NGC 4486B with $M_{\rm BH}=3.6\times 10^{8}\,M_{\odot}$ and $R_{c}\sim 6$ pc (taking the double nucleus separation of 12 pc to be the distance between the periapsis and apoapsis of the eccentric disk), we estimate (using eq. 2), a required recoil kick $V_{\rm{kick}}\approx 340$$\mathrm{km\ s}^{-1}$. This is a fairly moderate magnitude for a GW recoil kick (see Figure 4).
In Figure 4, we show the expected recoil kick magnitude as a function of the pre-merger black hole mass ratio, $q=m_{2}/M_{1}$ (where $M_{1}$ is the more massive black hole in the binary); adapted from Akiba.2025 applying the analytical prescription given in Lousto.2010; Lousto.2012; Fragione.2023. The light purple region shows the Monte Carlo results of $10^{6}$ evaluations of the analytical model (described in Fragione.2023) with randomized spins. The dimensionless spins for both black holes are drawn from an isotropic distribution in direction and a uniform distribution in magnitude ([0, 1]). The thick purple line shows the average kick magnitude as a function of $q$ marginalized over the spin distribution. The mean and $1\sigma$-confidence interval of the recoil kick distribution is shown, and the estimated $V_{\mathrm{kick}}\sim 340\,\mathrm{km\ s}^{-1}$ , required to induce NGC 4486B’s double nucleus is indicated with the dashed gray line. This recoil kick implies a pre-merger mass ratio of $q\approx 0.24$ which corresponds to a predicted companion mass of $7\times 10^{7}\,M_{\odot}$ (the calculation does not account for mass lost to GWs). However, the estimated recoil kick magnitude falls within one standard deviation of the mean for all mass ratios $q>0.15$, so one can only determine a lower limit at the companion mass of $4.7\times 10^{7}\,M_{\odot}$. Additionally, it is important to note that these recoil kick and mass ratio estimates assume that stars are initially on circular orbits. Akiba.2023 showed that the inclusion of initial eccentricities reduces the kick magnitude required to produce an eccentric disk at a given characteristic radius, so a degeneracy exists between the companion mass estimate and the pre-merger stellar eccentricities in the nuclear star cluster. This analytic framework establishes the connection between the eccentric disk geometry and the GW recoil kick. We emphasize that the purpose of this analytic treatment is to assess the plausibility of this scenario. More careful calculations and simulations will be carried out in the future. We build upon the above analysis in the following sections to explore the plausible timescales of such an event and the dynamical signatures it may have imprinted on the double nucleus in NGC 4486B.
Figure 5: Trajectories of the black hole that is kicked from the center of NGC 4486B. The different colored curves represent kicks in different directions. The top panels show the BH’s relative distance from the galaxy center, and the bottom panels show its relative speed. The left and right panels correspond to a kick velocity of $V_{\rm kick}=0.2V_{\rm esc}$, $0.5V_{\rm esc}$, and $0.75V_{\rm esc}$, respectively. The black dashed and dotted lines show the expected amplitude of Brownian motion and the observed offset in position and $v_{\mathrm{los}}$, respectively. In all cases, the black hole quickly returns close to the galaxy center, with position and velocity offsets that are well below the observed values.
4 Post kick trajectory
Considering the strong evidence in support of a recoil kick, we now explore the possible post-kick trajectories of the SMBH. We construct an $N$-body representation of the best-fit Schwarzchild model of NGC 4486B with $10^{6}$ particles. This model, which has no dark matter halo, has a best-fit black hole mass of $M_{\rm BH}=3.51\times 10^{8}\mbox{$M_{\odot}$}$, total stellar mass $M_{*}=7.44\times 10^{9}\mbox{$M_{\odot}$}$, and mass-to-light ratio $\Upsilon_{[F850LP]}=4.1$. We introduce a black hole particle at the center of the system and give it a kick velocity $V_{\rm kick}$ along several directions. We explore $V_{\rm kick}=0.2V_{\rm esc}=154\,\mathrm{km\ s}^{-1}$, $V_{\rm kick}=0.5V_{\rm esc}=384\,\mathrm{km\ s}^{-1}$, and $V_{\rm kick}=0.75v_{\rm esc}=576\,\mathrm{km\ s}^{-1}$, where $V_{\rm esc}$ is the escape velocity from the galaxy center (the middle value being close to the kick velocity estimated in Section 3.2).
The simulations are run using the publicly available code GyrfalcOn (Dehnen2002). Each particle is softened using the P1 softening kernel [^1] A modification of the Plummer softening kernel whose density is $\propto(1+x^{2})^{-7/2}$, with softening length 0.79 pc for the stars and 0.395 pc for the BH. We use hierarchical timesteps, where the timestep of each particle depends on it softening length $\epsilon$ and instantaneous acceleration $a_{i}$ as $\Delta t\sim 0.025\sqrt{\epsilon/\left|a_{i}\right|}$. According to this criterion, a circular orbit around the black hole at 0.79 pc (the softening length) is resolved with at least $\sim$100 timesteps.
Figure 5 shows properties of the resulting trajectory of the SMBH in the various cases. The top and bottom panels show the SMBH’s distance and velocity with respect to the galaxy’s center of mass [^2] In order to calculate the center of mass, we use the $90\%$ most bound stars in the simulation, as some stars may get unbound during the recoil kick. The results differ only slightly based on the exact center of mass definition., and the right, middle, and left panels correspond to cases with $V_{\rm kick}=0.2V_{\rm esc}$, $0.5V_{\rm esc}$, and $0.75V_{\rm esc}$, respectively. Notably, in all simulations, the line-of-sight velocity dispersion peak consistently follows the SMBH, justifying our assumption in Section 3.1 that the off-center velocity dispersion peak marks the location of the SMBH. Within each panel, the various curves correspond to different initial kick directions. In all panels, the dotted black line marks the observed positional offset of the SMBH (6 pc) and the localized $v_{\mathrm{los}}$ offset (16 km s-1) measured at the location of the SMBH (see top panel of Fig. 2). This feature is assumed to trace the mean motions of stars bound to the SMBH but may not represent the intrinsic 3D velocity of the black hole itself. However, for comparison with the simulations, we refer to this quantity as the “observed line-of-sight velocity”. The black dashed lines in the lower and upper panels show the expected velocity and position amplitudes due to Brownian motion, respectively, which are given by Gualandris.2008:
$V_{\rm Brown}=\sigma_{0}\sqrt{\frac{3m_{*}}{M_{\rm BH}}}\quad\quad R_{\rm Brown}=\frac{V_{\rm Brown}}{\sqrt{4\pi G\rho_{c}/3}}$ (3)
where $\sigma_{0}$ is the central velocity dispersion, $m_{*}$ is the mass of each star particle, and $\rho_{c}$ is the central density.
In every trajectory shown in Figure 5, the black hole rapidly sinks back toward the galaxy center, reaching amplitudes comparable to the observed positional and $v_{\mathrm{los}}$ offsets, after which these quantities fall below the observed values. The return timescale increases with kick strength, from $\sim$10 Myr for $0.2V_{\rm esc}$ to $\sim$30 Myr for $0.5V_{\rm esc}$ and $\sim$80 Myr for $0.75V_{\rm esc}$. Although these timescales differ, all are extremely short relative to a Hubble time. Furthermore, the observed $v_{\mathrm{los}}$ offset is only one component of the possible 3D velocity of the SMBH. At late times, the black hole retains a residual velocity relative to the galaxy center that is clearly resolved above the expected Brownian amplitude. This residual motion is likely the result of a core-stalling effect (e.g., Read2006), but it remains insufficient to account for the observed velocity offset. Interestingly, the timescale for the black hole to return to the center is longer if the kick is along the Z-axis, compared to the other directions, although this timescale is still significantly shorter than a Hubble time. This will be further investigated in a follow-up paper (Jha et al, in preparation).
[…]
It must be noted that the initial galaxy models we used for the simulations are based on the current observations of NGC 4486B. As discussed previously, oscillations of the SMBH following a recoil kick can scatter stars and flatten the central cusp beyond what would have resulted from “binary black hole scouring” alone. The density profile of NGC 4486B prior to the kick was therefore likely steeper (denser) than assumed in our simulations. A denser central region would enhance dynamical friction and further shorten the return timescale, implying that the black hole return times inferred from the trajectories in Figure 5 likely represent upper limits. In a forthcoming study, we will extend this analysis by modeling recoil trajectories in more strongly cusped pre-kick galaxy profiles.
In the following section, we investigate additional dynamical scenarios—such as buoyancy-driven oscillations and the stalling of binary SMBHs in rotating systems—to evaluate whether they can account for the observed photometric and kinematic features. These comparisons allow us to assess whether a recent recoil remains the most compelling explanation, or whether alternative processes could account for the observed off-centered nucleus.
5 Alternative Dynamical Scenarios for Black Hole Displacement
5.1 Buoyancy-Induced Black Hole Offset
Dynamical buoyancy is the opposite of dynamical friction, where the host system exerts a net positive (outward) torque on a massive perturber, leading to its outward motion (Cole2012; Banik2021; Banik2022; Dattathri2025b). This phenomenon is typically associated with cored density profiles. Since NGC 4486B has a central core (Figure 1), we investigate here whether dynamical buoyancy operates, and if so, what the associated offsets are.
We initialize an $N$-body simulation with the same setup and parameters as Section 4, but without any recoil kick, i.e., the SMBH is initially at rest at the galaxy center. We then let the system evolve and track the SMBH’s motion around the galaxy center. Figure 6 shows the resulting trajectory of the SMBH’s position (top) and velocity (bottom). Similar to Figure 5, the black dashed line shows the expected Brownian motion amplitude, and the black dotted line shows the observed offsets. The SMBH shows a clear initial outward motion even without a recoil kick. This motion is well resolved above the Brownian amplitude, indicating that it is not a numerical artifact. The offset reaches a maximum amplitude of $\sim 50\,\mathrm{km\ s}^{-1}$ and $4$ pc away from the center at $T\sim 2$ Myr. However, similar to the recoil simulations (Figure 5), the black hole quickly sinks back to the center. After $T\gtrsim 15$ Myr, the SMBH is orbiting around the galaxy center at a speed of $\sim 5$$\mathrm{km\ s}^{-1}$and a distance 121-2 pc, which is significantly lower than the observed offsets.
The origin of the buoyancy in our system is unclear. Dattathri2025b suggested that in spherical isotropic systems, dynamical buoyancy is due to an unstable dipole mode in the system. This instability may arise if the distribution function has a region where ${\rm d}f/{\rm d}E>0$ (Dattathri2025a). However, NGC 4486B is not isotropic and has a net rotation (Figure 2); hence, it is unclear whether the same instability would occur here. A recent study by Boily2025 analyzed the motion of BHs in anisotropic galaxy cores and found that if the system has an excess of streaming motion over its velocity dispersion, a central black hole can gain angular momentum and move outward. Both of the above processes could work in tandem to produce an offset SMBH. However, regardless of the mechanism, it is clear from Figure 6 that if the SMBH in NGC 4486B was initially at rest in the center, it could attain a significant velocity offset due to dynamical buoyancy for a short period of time $\sim 6-7$Myr, but the position offset is always well below the observed value of 6 pc. Note that the observed 6 pc separation is a projected distance; the true 3D offset can only be larger, which further strengthens this argument.
Figure 6: Trajectory of the black hole that is placed at the center of NGC 4486B without any recoil kick. The black hole shows a clear outward motion away from the galactic center that is well resolved above the Brownian amplitude. However, except for very early times, the associated offsets are well below the observed values.
5.2 Pre-Merger Binary Black Hole in Prograde Orbit
Figure 7: Time evolution of the galaxy and the SMBH binary parameters. Top left: Volume-density profiles of the NGC 4486B model at different times during the SMBH binary evolution. Top right: Orbit of the SMBH binary in the reference frame of the stellar density center. The time-colored line shows the trajectory of the binary CoM after subtracting the density center motion; the density center itself is shown at the origin with a red dot. Bottom left: Velocity of the SMBH binary CoM relative to the density center, illustrating the residual motion of the binary within the host galaxy. Bottom right: Velocities of the individual SMBHs relative to the binary CoM, showing that the lighter secondary SMBH moves faster than the primary and that both velocities increase as the binary hardens.
Although NGC 4486B is overall a dispersion-dominated galaxy, it exhibits a considerable central rotation ($|v|/\sigma\simeq$ 0.40-0.65). Studies have shown that even this level of rotational support can profoundly influence the dynamics of SMBH binaries in rotating hosts (holley+15; Rasskazov+16; Mirza+17; Khan2020). In particular, simulations demonstrate that when a binary is on a prograde orbit aligned with the stellar rotation, its center-of-mass (CoM) can settle into a stable orbit around the galactic center with a characteristic radius comparable to the influence radius of the two BHs. Once established, such a system can survive for extended periods, maintaining a corotating orbit while the SMBHs in binary continue to shrink their relative separation, ultimately coalescing following GW emission.
This scenario is especially relevant for NGC 4486B. Photometric and kinematic studies indicate that it is an overall old, dynamically relaxed galaxy: its stellar populations are predominantly ancient, its large-scale light distribution is smooth, and no evidence exists for recent large-scale tidal disturbances (e.g., Ferrarese.2006; Guerou.2015; Janz.2016). Thus, any major merger that contributed to its present structure likely occurred several gigayears ago (Khan2018; Kelly2025), consistent with its environment deep within the Virgo cluster. Nevertheless, if that merger brought in a secondary SMBH on a prograde trajectory, the resulting binary could have settled into a long-lived corotating orbit before the SMBHs achieved a coalescence. This outcome naturally extends the effective lifetime of the binary off-center phase and is supported by the $N$-body simulations of rotating systems cited above.
[…]
Once again, our initial $N$-body model reflects the current structure of NGC 4486B rather than that of its progenitor, which likely possessed a steeper central density cusp prior to binary scouring. Even so, the simulation demonstrates that in a host galaxy with the modest but non-negligible large-scale rotation observed, a forming SMBH binary can become trapped in a corotation resonance for a finite period. Because of the high computational cost of direct $N$-body integrations, we have evolved this model only to $\sim 30$ Myr. Nevertheless, our results indicate that while NGC 4486B is globally an old and relaxed system, its central SMBH binary could plausibly have remained in a long-lived resonant, corotating configuration following a galaxy–galaxy merger in the cluster outskirts (the Virgo cluster crossing time is $\sim 10^{9}$ yr), and may have undergone coalescence only relatively recently. Based on our estimates, the SMBH binary would require $\approx 500~\mathrm{Myr}$ to coalesce after forming via dynamical-friction–driven inspiral—an inspiral that itself may take several hundred Myr if the secondary SMBH arrives at a separation of $\sim 1~\mathrm{kpc}$ during the preceding merger.
In a forthcoming study, we will explore this scenario using a larger suite of long-duration direct $N$-body simulations—initiated from a more strongly cusped pre-scouring density profile—and follow the evolution through GW–driven coalescence, with the goal of determining whether the merged SMBH can retain a sustained orbit around the galaxy’s density center.
6 Summary and Conclusion
Although the double nucleus in NGC 4486B was first discovered nearly 30 years ago (Lauer.1996) there have been few studies of its origin and dynamical structure. In analogy with the double nucleus in M31, Lauer.1996 proposed that it was an apsidally aligned eccentric nuclear disk surrounding an SMBH. Two striking differences between M31 and NGC 4486B however are: (a) in M31 the SMBH (located at P3) is at the galactic center while the SMBH in NGC 4486B is 6pc away from the center (defined by the large scale isophotes) and possibly has a velocity offset of $\sim 16\,\mathrm{km\ s}^{-1}$ relative to stars on the opposite side of the galaxy; (b) M31 has a central stellar cusp and a relatively young star cluster coincident with the SMBH while NGC 4486B has a flat core $\gamma=0.14$ with a core radius of $\sim 20$ pc. In Paper 1 (Behzad.2025) we used JWST-NIRSpec IFU spectroscopy and archival HST photometry combined with orbit-superposition Schwarzschild dynamical modeling to confirm the presence of an SMBH of ${\cal M}_{BH}=3.6\pm 10^{8}\mbox{$M_{\odot}$}$. Despite the fact that the Schwarzschild model is axisymmetric and consequently does not reproduce the asymmetric photometry and kinematics, the distribution function provided by the model allows us to examine the origin of the double nucleus and the offset SMBH in NGC 4486B for the first time.
In this work, we focus on the possibility that the double nucleus in NGC 4486B is due to an END similar to that in M31 (Tremaine.1995). We compare the orbital properties of the Schwarzschild model with predicted properties of ENDs that are produced when an SMBH surrounded by a circular nuclear disk experiences a GW recoil kick. These simulations show that a GW recoil following SMBH coalescence can transform an initially circular stellar disk into an eccentric, apsidally aligned structure with a significant retrograde population. Previous studies of surface brightness and kinematic distributions of simulated ENDs match many properties of the observed double nucleus in NGC 4486B: the fainter peak and the velocity dispersion peak coincide and are the location of the SMBH, the asymmetric line-of-sight velocity field, are all consistent with predictions for nearly edge-on ENDs. This interpretation is further reinforced by the finding that our best-fitting Schwarzschild models require a significant fraction of retrograde orbits in the outer nuclear region. Retrograde orbits are not only a natural dynamical consequence of a recoiling SMBH but also a key ingredient in the long-term stability of ENDs. The presence of the central core in this galaxy is additional evidence that it has experienced a binary-black hole mediated “cusp scouring” prior to the binary black hole merger that likely gave the resulting SMBH the kick which produced the END.
From the observed properties of the double nucleus (semi-major axis $\sim 6$ pc) and black hole mass we estimate that the SMBH experienced a recoil kick of $\sim 340\,\mathrm{km\ s}^{-1}$. We adapt the formalism of (Akiba.2025) to estimate the recoil kick magnitude as a function of the pre-merger black hole mass ratio, $q=m_{2}/M_{1}$. Marginalizing over the spin distribution of the two BHs (which is assumed to be isotropic in direction and uniform in scaled magnitude between [0,1]), we estimate that for a recoil kick of $340\,\mathrm{km\ s}^{-1}$, the pre-merger mass ratio was: $q\sim 0.24\pm 0.09$ implying that the mass of the secondary was $m_{2}\sim 7\pm 2.3\times 10^{7}\mbox{$M_{\odot}$}$.
To explore this scenario further, we constructed $N$-body realizations of our best-fit Schwarzschild model from Paper 1 and carried out numerous experiments, which show that a recoiling SMBH, kicked with a velocity of $0.2-0.75v_{\rm esc}\sim 154-576\,\mathrm{km\ s}^{-1}$, oscillates and sinks back to the galactic center within $\sim$10-80 Myr. For our estimated recoil kick of $340\,\mathrm{km\ s}^{-1}$, the timescale to sink to the center is expected to be less than 30 Myr, implying that if the position and velocity offsets of the SMBH are to be attributed to a GW recoil kick, the binary black hole merger in NGC 4486B must have occurred quite recently. The simulations rule out “core stalling” and “dynamical buoyancy” scenarios in which the failure of dynamical friction in a constant density core would allow the SMBH to oscillate for extended periods (Dattathri2025b). This discrepancy is likely due to the overmassive nature of the SMBH in NGC 4486B (${\cal M}_{BH}/M_{*}\sim 4-10$%), while most numerical simulations of this problem assume ${\cal M}_{BH}/M_{*}\sim 0.1-0.2$%, based on expectations of SMBH scaling relations.
Our inference that the recoil event occurred recently presents an apparent paradox. NGC 4486B lies near the dense center of the Virgo Cluster and exhibits a uniformly old, dynamically relaxed stellar population. The absence of tidal features and the uniformly old stars imply that the $\sim 1:4-1:5$ major merger that introduced the secondary black hole must have taken place several gigayears ago, likely when the progenitor galaxies were still in the cluster outskirts.
[…]
The implications of this finding extend beyond a single galaxy. If compact ellipticals with double nuclei arising from ENDs like NGC 4486B retain signatures of relatively recent SMBH mergers, they may provide valuable laboratories for studying post-merger dynamics on resolvable scales. Non-equilibrium dynamical models similar to those constructed for M31 (Peiris.2003; Bacon.2001; Sambhus.2002; BrownMagorrian.2013) fitted to the excellent NIRSpec-IFU kinematic maps of NGC 4486B and $N$-body models of the resulting END, embedded inside the rotating elliptical host are needed to more precisely determine the time since the merger event and the expected survival time of the END. NGC 4486B serves not only as a rare nearby case study, but also as a cornerstone for linking high-resolution observations of galactic centers to GW astrophysics.
We are grateful to Tod Lauer for kindly providing the deconvolved F [^5] W image from (Lauer.1996). We thank Eugene Vasiliev for useful discussions. BT and MV acknowledge funding from Space Telescope Science Institute awards: JWST-GO-02567.002-A and HST-GO-16882.002-A. M.A.T. and S.T. acknowledge the support of the Canadian Space Agency (CSA) [22JWGO1-07]. This work was supported by Tamkeen under the NYU Abu Dhabi Research Institute grant CASS, and utilized the high-performance computing resources at New York University Abu Dhabi. CL acknowledges support from the National Natural Science Foundation of China (NSFC, Grant No. 12173025), National Key R&D Program of China (2023YFA1607800, 2023YFA1607804), 111 project (No. B20019), and Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education. B.T. acknowledges support from the Mendel Science Experience (MSE) Postdoctoral Fellowship awarded by Villanova University.
Figure A1: Best-fit Schwarzschild models with (top) and without (bottom) retrograde orbits, each annotated with the corresponding $\chi^{2}$ values.